OpenStudy (anonymous):

-2/5 x - 9 < 9/10???

OpenStudy (anonymous):

Can you help me???

OpenStudy (anonymous):

i might probably get this wrong (i have so many people here to correct me :D) q+12-2(q-22)> 10 2(q-22)+q+12>10+2(q-22) additive inverse q+12> 10+2(q-22) q+22>10+ 2q-44 simplify (34)+ q+22 > 2q-34 +(34) additive inverse q+ 56 > 2q 56> q

OpenStudy (anonymous):

Umm idk... Can anyone else give there opinion?

OpenStudy (paxpolaris):

\[-\frac25 x - 9 < \frac9{10}\]First You want to isolate the monomial with x, so:\[\implies -\frac25 x<\frac9{10}+9\] follow so far?

OpenStudy (paxpolaris):

\[9+ \frac9{10}= {99\over10}\]\[\therefore -\frac25 x<{99 \over 10}\] Next you want to get rid of the co-efficient of x...

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

lol wrong problem

OpenStudy (anonymous):

that was for the last post

OpenStudy (anonymous):

i copy and pasted

OpenStudy (paxpolaris):

You do that by multiplying both sides by -5/2... Note that when multiplying both sides by a negative number flips the < sign to >\[-\frac52 \cdot-\frac25\cdot x>-\frac52\cdot {99 \over 10}\]\[\implies 1\cdot x>-{99 \over 4}\]\[\implies x>-24{3 \over 4}\]

OpenStudy (anonymous):

thanks now can u answer this? Starting from 1.5 miles away, a car drives towards a speed checkpoint and then passes it. The car travels at a constant rate of 53 miles per hour. The distance of the car from the checkpoint is given by . At what times is the car 0.1 miles from the checkpoint? Calculate your answer in seconds